OpenStudy (anonymous):

what is the numerical value of the following 2^6

OpenStudy (anonymous):

2^6 = 64 64 + 1 = 65

OpenStudy (anonymous):

\[2^6 = 64.\]

OpenStudy (anonymous):

In binary?